# On the Geometry of Deep Learning

### v1

@@ Introduction
Interpolation is the point that would typically be recommended to stop training and fix the network for use in an application. But let’s see what happens if we continue training about 37 times longer. At right in Figure 9, we see that, while the training error has not changed after continued training (it is still near zero, meaning correct classification of nearly all the training data), the tessellation has metamorphosed. There are now only half as many tiles in this region, and they have all migrated to define the decision boundary, where presumably they are being used to create sharp decisions. Around the training data, we now have a very low density of tiles with low 2-norm of their $\boldsymbol{A}_{\omega}$ matrices, and thus presumably a much smoother mapping $f(\boldsymbol{x})$. Hence, the sensitivity of $f(\boldsymbol{x})$ as measured by a local Lipshitz constant will be much lower than just after interpolation.
@@
We designate this state of affairs delayed robustness; it is one facet of the general phenomenon of grokking that has been discovered only recently [power2022grokking]. A dirty secret of deep networks is that $f(\boldsymbol{x})$ can be quite unstable to small changes in $\boldsymbol{x}$ (which seems expected given the high degree of nonlinearity). This instability makes deep networks less robust and more prone to attacks like causing a ‘barn’ image to be classified as a ‘pig’ by adding a nearly undetectable but carefully designed attack signal to the picture of a barn. Continuing learning to achieve grokking and delayed robustness is a new approach to mitigating such attacks in particular and making deep learning more stable and predictable in general.
@@
Can we translate the visualization of Figure 9 into a metric that can be put into practice to compare or improve deep networks? This is an open research question, but here are some first steps [alwaysgrok]. Define the local complexity (LC) as the number of tiles in a neighborhood $V$ around a point $\boldsymbol{x}$ in the input space. While exact computation of the LC is combinatorially complex, an upper bound can be obtained in terms of the number of hyperplanes that intersect $V$ according to Zaslavsky’s Theorem, with the assumption that $V$ is small enough that the hyperplanes are not folded inside $V$. Therefore, we can use the number of hyperplanes intersecting $V$ as a proxy for the number of tiles in $V$.
@@
For the experiment reported in Figure 9, we computed the LC in the neighborhood of each training data point in the entire training dataset and then averaged those values. From the above discussion, high LC around a point $\boldsymbol{x}$ in the input space implies small, dense tiles in that region and a potentially unsmooth and unstable mapping $f(\boldsymbol{x})$ around $\boldsymbol{x}$. The values reported in Figure 9 confirm that the LC does indeed capture the intuition that we garnered visually. One interesting potential application of the LC is as a new progress measure that serves as a proxy for a deep network’s expressivity; LC is task-agnostic yet informative of the training dynamics.
@@
Open research questions regarding the dynamics of deep network learning abound. At a high level, it is clear from Figure 9 that the classification function being learned has its curvature concentrated at the decision boundary. Approximation theory would suggest that a free-form spline should indeed concentrate its tiles around the decision boundary to minimize the approximation error. However, it is not clear why that migration occurs so late in the training process.
@@
Another interesting research direction is the interplay between grokking and batch normalization, which we discussed above. Batch normalization provably concentrates the tessellation near the data samples, but to grok we need the tiles to move away from the samples. Hence, it is clear that batch normalization and grokking compete with each other. How to get the best of both worlds at both ends of the gradient learning timeline is an open question.
@@
Deep generative models map from typically a low-dimensional Euclidean input space (called the parameter space) to a manifold ${\cal M}$ of roughly the same dimension in a high-dimensional output space. Each point $\boldsymbol{x}$ in the parameter space synthesizes a corresponding output point $\widehat{\boldsymbol{y}}=f(\boldsymbol{x})$ on the manifold (e.g., a picture of a bedroom). Training on a large number of images $\boldsymbol{y}_{i}$ learns an approximation to the mapping $f$ from the parameter space to the manifold. It is beyond the scope of this review, but learning the parameters of a deep generative model is usually more involved than simple gradient descent [goodfellow2016deep]. It is useful for both training and synthesis to view the points $\boldsymbol{x}$ from the parameter space as governed by some probability distribution, e.g., uniform over a bounded region of the input space.
In the case of a GAN based on ReLU or similar activation functions, the manifold ${\cal M}$ is a continuous piecewise affine manifold; [^3] We allow ${\cal M}$ to intersect itself transversally in this setting. see Figure 10. Points on the manifold are given by (5) as $\boldsymbol{x}$ sweeps through the input space.
@@
A major issue with deep generative models is that, if the training data is not carefully sourced and currated, then they can produce biased outputs. A deep generative model like a GAN or VAE is trained to approximate both the structure of the true data manifold from which the training data was sampled and the data distribution on that manifold. However, all too often in practice, training data are obtained based on preferences, costs, or convenience factors that produce artifacts in the training data distribution on the manifold. Indeed, it is common in practice for there to be more training data points in one part of the manifold than another. For example, a large fraction of the faces in the CelebA dataset are smiling, and a large fraction of those in the FFHQ dataset are female with dark hair. When one samples uniformly from a model trained with such biased data, the biases will be reproduced when sampling from the trained model, which has far-reaching implications for algorithmic fairness and beyond.
We can both understand and ameliorate sampling biases in deep generative models by again leveraging their affine spline nature. The key insight for the bias issue is that the tessellation of the input space is carried over onto the manifold. Each convex tile $\omega$ in the input space is mapped to a convex tile $M(\omega)$ on the manifold using the affine transform
@@
In particular, it is easy to show that the mapping (9) from the input space to the manifold warps the tiles in the input space tessellation by $\boldsymbol{A}_{\omega}$, causing their volume to expand or contract by
@@
We can view deep networks as a tractable mechanism for emulating those most powerful of splines, the free-knot splines (splines like those in Figure 2 where the intervals partitioning the real line are arbitrary) in high dimensions. A deep network uses the power of a hyperplane arrangement to tractably create a myriad of flexible convex polytopal tiles that tessellate its input space plus affine transformations on each that result in quite powerful approximation capabilities in theory [devore2021neural] and in practice. There is much work to do in studying these approximations (e.g., developing realistic function approximation classes and proving approximation rates) as well as developing new deep network architectures that attain improved rates and robustness.
@@
We hope that we have convinced you that viewing deep networks as affine splines provides a powerful geomeric toolbox to better understand how they learn, how they operate, and how they can be improved in a principled fashion. But splines are just one interesting research direction in the mathematics of deep learning. These are early days, and there are many more open than closed research questions.

### v2

@@ Introduction
Interpolation is the point where most deep learning practitioners would stop training and fix the network’s weights and biases for use in their target application. But let’s see what happens if we continue training about 37 times longer. At right in Figure 9, we see that, while the training error does not improve after continued training (it is still near zero, meaning correct classification of nearly all the training data), the tessellation has metamorphosed. There are now only half as many tiles in this region, and they have nearly all migrated to define the decision boundary, where presumably they are being used to create sharp decisions. Around the training data, we have an extremely low density of tiles with low 2-norm of their $\boldsymbol{A}_{\omega}$ matrices and thus presumably a much smoother mapping $f(\boldsymbol{x})$. Hence, the sensitivity of $f(\boldsymbol{x})$ around the training data as measured by a local Lipshitz constant will be much lower than just after interpolation.
@@
This situation is an example of grokking, in which some desired property of a deep network unexpectedly occurs well after the training error converges to near zero [power2022grokking]. In this case, the desired property is the robustness of the network to perturbations in the training data. A dirty secret of today’s deep networks is that $f(\boldsymbol{x})$ can be quite unstable to small changes in $\boldsymbol{x}$ (which seems expected given their high degree of nonlinearity). This instability makes deep networks less robust and more prone to so-called adversarial attacks such as causing a ‘barn’ image to be classified as a ‘pig’ by adding a nearly undetectable but carefully designed attack signal to an image of a barn. Continuing learning to achieve grokking and delayed robustness is a new approach to mitigating such attacks in particular and making deep learning more stable and predictable in general.
@@
Can we translate the visualization of Figure 9 into a metric that can be put into practice to compare or improve deep networks? This is largely an open research question, but here are some first steps [alwaysgrok]. Define the local complexity (LC) as the number of tiles in a neighborhood $V$ around a point $\boldsymbol{x}$ in the input space. While exact computation of the LC is combinatorially complex, an upper bound can be obtained in terms of the number of hyperplanes that intersect $V$ according to Zaslavsky’s Theorem, with the assumption that $V$ is small enough that the hyperplanes are not folded inside $V$. Therefore, we can use the number of hyperplanes intersecting $V$ as a proxy for the number of tiles in $V$.
@@
For the experiment reported in Figure 9, we computed the LC in the neighborhood of each data point in the entire training dataset and then averaged those values. From the above discussion, high LC around a point $\boldsymbol{x}$ in the input space implies a multitude of small tiles in that region and a potentially unsmooth and unstable mapping $f(\boldsymbol{x})$ around $\boldsymbol{x}$. The values reported in Figure 9 confirm that the LC does indeed capture the intuition that we garnered visually.
@@
Open research questions regarding the dynamics of deep network learning abound. At a high level, it is clear from Figure 9 that the classification function ultimately being learned has its curvature concentrated at the decision boundary. Approximation theory would agree that a free-form spline should indeed concentrate its tiles around the decision boundary to minimize the approximation error. However, it is not clear why the migration occurs so late in the training process.
Another interesting research direction involves incorporating the average LC around the training data points in the optimization cost function (e.g., (4)) in order to encourage the network to converge to a stable mapping $f(\boldsymbol{x})$ sooner rather than much later.
@@
Yet another interesting research direction is the interplay between grokking and batch normalization, which we discussed above. Batch normalization provably concentrates the tessellation near the training data points, but (at least for classification problems) to grok we need the tiles to move away from those points towards the decision boundaries. Hence, it is clear that batch normalization and grokking compete with each other. How to get the best of both worlds at both ends of the gradient learning timeline is an open question.
@@
Deep generative models typically map from a low-dimensional Euclidean input space (called the parameter space) to a manifold ${\cal M}$ of roughly the same dimension in a high-dimensional output space. Each point $\boldsymbol{x}$ in the parameter space synthesizes a corresponding output point $\widehat{\boldsymbol{y}}=f(\boldsymbol{x})$ on the manifold (e.g., a picture of a bedroom). Training on a large number of data points $\boldsymbol{y}_{i}$ learns an approximation to the mapping $f$ from the parameter space to the manifold. It is beyond the scope of this review, but learning the parameters of a deep generative model is usually more involved than simple gradient descent [goodfellow2016deep]. It can be useful for both training and synthesis to view the points $\boldsymbol{x}$ from the parameter space as governed by some probability distribution, e.g., uniform over a bounded region of the input space.
For a deep generative model based on a ReLU or similar activation function, the manifold ${\cal M}$ is a continuous piecewise affine manifold; [^3] We allow ${\cal M}$ to intersect itself transversally in this setting. see Figure 10. Points on the manifold $f(\boldsymbol{x})$ are given by (5) as the parameter vector $\boldsymbol{x}$ sweeps through the input parameter space.
@@
A major issue with deep generative models is that, if the training data is not carefully sourced and currated, then they can produce biased outputs. A deep generative model like a GAN or VAE is trained to approximate both the structure of the true data manifold from which the training data was sampled and the data distribution on that manifold. However, all too often in practice, training data are obtained based on preferences, costs, or convenience factors that produce artifacts in the training data distribution on the manifold. Indeed, it is common in practice for there to be more training data points in one part of the manifold than another. For example, a large fraction of the faces in the CelebA dataset are smiling, and a large fraction of those in the FFHQ dataset are female with dark hair. When one samples uniformly from a model trained using such biased data, the biases are reproduced when sampling from the trained model, which has far-reaching implications for algorithmic fairness and beyond.
We can both understand and ameliorate sampling biases in deep generative models by again leveraging their affine spline nature. The key insight for the bias issue is that the tessellation of the input parameter space is carried over onto the manifold. That is, each convex tile $\omega$ in the input space is mapped to a convex tile $M(\omega)$ on the manifold using the affine transform
@@
In particular, it is easy to show that the mapping (9) from the input space to the manifold shears the tiles in the input space tessellation by $\boldsymbol{A}_{\omega}$, causing their volume to expand or contract by the factor
@@
We can view deep networks as a tractable mechanism for emulating those most powerful of splines, the free-knot splines (splines like those in Figure 2 where the intervals partitioning the real line domain are arbitrary) in high dimensions. A deep network uses the power of a hyperplane arrangement to tractably create a myriad of flexible convex polytopal tiles that tessellate its input space plus affine transformations on each that result in quite powerful approximation capabilities in theory [devore2021neural] and in practice. There is much work to do in studying these approximations (e.g., developing realistic function approximation classes and proving approximation rates) as well as developing new deep network architectures that attain improved rates and robustness.
@@
We hope that we have convinced you that viewing deep networks as affine splines provides a powerful geomeric toolbox to better understand how they learn, how they predict, and how they can be improved in a principled fashion. But splines are just one interesting research direction in the mathematics of deep learning. These are early days, and there are many more open than closed research questions.